# A.S.E: A Repository-Level Benchmark for Evaluating Security in AI-Generated Code

### v2

@@ Abstract
The increasing adoption of large language models (LLMs) in software engineering necessitates rigorous security evaluation of their generated code. However, existing benchmarks often lack relevance to real-world AI programming scenarios, making them inadequate for assessing the practical security risks associated with AI-generated code in production environments. To address this gap, we introduce A.S.E (AI Code Generation Security Evaluation), a repository-level evaluation benchmark designed to closely mirror real-world AI programming tasks, offering a comprehensive and reliable framework for assessing the security of AI-generated code. Our evaluation of leading LLMs on A.S.E reveals several key findings. In particular, current LLMs still struggle with secure coding. The complexity in repository-level scenarios presents challenges for LLMs that typically perform well on snippet-level tasks. Morever, a larger reasoning budget does not necessarily lead to better code generation. These observations offer valuable insights into the current state of AI code generation, assisting developers in selecting the most appropriate models for practical tasks, while laying the foundation for refining LLMs to generate secure and efficient code in real-world applications.
@@ 1 Introduction
The rapid advancement of large language models (LLMs) has greatly enhanced the AI programming ecosystem, with tools like Cursor and Claude Code enabling developers to choose models that best fit their tasks. These AI assistants significantly improve programming efficiency, leading to a surge of AI-generated code in production environments. However, research @8SecurityEval ; DBLP:journals/corr/abs-2502-11844@12BaxBench ; DBLP:conf/llm4code/PengCHYR25@11CWEval ; DBLP:conf/satml/HajipourHHSF24@9CodeLMSec ; DBLP:journals/corr/abs-2506-05692@13SafeGenBench ; fu2024constraineddecodingsecurecode has shown that such code can harbor security vulnerabilities, posing serious risks such as data breaches or system failures. Relying on developers to ensure the security of AI-generated code can be highly challenging given the complexity of modern software systems. Therefore, there is a pressing need to identify and utilize AI models that are capable of generating secure code.
@@
Despite the numerous benchmarks developed by both academia and industry to evaluate AI-generated code, most of them  DBLP:journals/corr/abs-2107-03374@17HumanEval ; DBLP:journals/corr/abs-2108-07732@19MBPP ; DBLP:conf/iclr/0003XM24@28RepoBench ; DBLP:journals/corr/abs-2406-11612@31Long-code-arena ; DBLP:conf/nips/DingWADTJRNBRX23@29CrossCodeEval ; DBLP:conf/acl/LiangJH025@32REPOCOD ; DBLP:conf/acl/LiZGM0PHWL25@33FEA-Bench ; jimenez2024swebench primarily focus on code quality, such as syntax correctness and functional accuracy, while overlooking critical security considerations. Although some benchmarks @8SecurityEval ; DBLP:journals/corr/abs-2502-11844@12BaxBench ; DBLP:conf/llm4code/PengCHYR25@11CWEval ; DBLP:conf/satml/HajipourHHSF24@9CodeLMSec ; DBLP:journals/corr/abs-2506-05692@13SafeGenBench ; fu2024constraineddecodingsecurecode attempt to address code security (as shown in Table 1), they are often inadequate to assess the actual security risks of AI-generated code in real-world production scenarios due to several key reasons: (a) Limited relevance to real-world data. These datasets are typically sourced from human-curated synthetic code snippets, which have limited relevance to the functional and security scenarios of real-world projects. (b) Code generation tasks detached from real-world AI programming. Their code generation tasks are typically limited to isolated code snippets, focusing solely on functional descriptions without considering the context within files or projects, which does not align with mainstream AI programming paradigm. (c) Unreliable code evaluation methods. Security assessments of generated code often rely on manual or LLM-based judgment, which are unreliable and difficult to automate or reproduce consistently. This gap poses a significant challenge for both developers and organizations seeking to integrate AI-generated code securely into their systems.
@@
Building on these design principles, the A.S.E benchmark includes 120 repository-level instances, consisting of 40 seed datasets collected from GitHub and 80 mutated variants generated through semantic and structural mutation techniques, such as identifier renaming and control-flow reshaping. These variants are introduced to mitigate data leakage risks, ensuring that the evaluation reflects the LLM’s capabilities rather than its memorization. To closely simulate real-world code generation processes, A.S.E uses the BM25 algorithm to automatically extract relevant code context from the repository when designing code generation tasks. These code generation tasks cover four common vulnerability types in real-world web projects. We focus on the web domain due to the primary application of AI programming in web development in production environments. These vulnerability-related codes to be generated are closely tied to the projects’ business logic, requiring models to understand both project code and security considerations. A.S.E also incorporates five commonly used web development languages to assess how programming languages influence model performance. Furthermore, A.S.E integrates customized SAST rules for security evaluation on each data instance in the dataset. While A.S.E aims to evaluate the security of AI-generated code, it also considers code quality and stability. Given that security is meaningful only when code correctness is ensured, and in light of potential hallucinations in large model outputs, A.S.E uses multi-dimensional metrics that assess code security, quality, and generation stability for a comprehensive evaluation of the model’s capabilities.
@@
Based on A.S.E, we evaluated 26 mainstream commercial and open-source models under the same experimental setup, leading to several key findings. First, existing LLMs still face significant challenges in secure coding. All models fall short in terms of security performance compared to their code quality performance (such as syntax correctness). Even the best-performing model, Claude-3.7-Sonnet, achieved only a total score of 52.79 in our evaluation. Second, A.S.E introduces significant complexity in repository-level scenarios, which presents a challenge for LLMs that typically perform well on snippet-level tasks. For example, although GPT-3 excels on SafeGenBench DBLP:journals/corr/abs-2506-05692@13SafeGenBench , its performance on the A.S.E. benchmark drops, falling behind many other models. Third, slow-thinking configurations, which allocate more deliberate computation or multi-step reflection, tend to underperform compared to fast-thinking configurations that rely on concise, direct decoding. This suggests that a larger reasoning budget does not necessarily lead to better code generation. These observations offer valuable insights into the current state of AI code generation, helping developers select the most appropriate models for their specific tasks. Furthermore, they provide a foundation for refining LLMs, enhancing their ability to generate secure and efficient code in real-world applications.
@@ 2 Related Work
Table 1: Comparison of Security-Oriented Code Generation Evaluation Datasets
@@
SafeGenBench Snippet Human-Curated Synthetic Task-Based Coding Problems SAST + LLM
@@
CWEval Snippet Human-Curated Synthetic Task-Based (CWE-Tagged) Test Cases
CODEGUARD+ Snippet Human-Curated Synthetic Task-Based Coding Problems Test Cases
CodeLMSec Snippet Human-Curated Synthetic Task-Based Coding Problems SAST
SecurityEval Snippet Human-Curated Synthetic Task-Based Coding Problems SAST + Manual
@@ 2.1 Relevance to real-world scenarios
2.2 Code Assessment Methods.
Prior security-oriented benchmarks have adopted diverse methods for assessing the security of AI-generated code, yet each comes with notable limitations. (a) Manual evaluation offers flexibility but is labor-intensive and difficult to scale. For example, SecurityEval @8SecurityEval relies on expert validation, which makes it difficult to automate and scale to new model evaluations. (b) LLM-as-judge approaches scale more easily and capture semantic nuances, but their judgments are highly sensitive to prompt design, model version, and decoding randomness, making results unreliable and hard to reproduce (DBLP:journals/corr/abs-2312-04724@25CyberSecEval, ). (c) Generic static analysis tools (SAST) provide deterministic rule-based detection, yet often suffer from false positives and false negatives when applied across different languages or without contextual calibration DBLP:conf/satml/HajipourHHSF24@9CodeLMSec ; DBLP:journals/corr/abs-2506-05692@13SafeGenBench . (d) Test-case-based assessments are straightforward but inherently limited DBLP:journals/corr/abs-2502-11844@12BaxBench ; DBLP:conf/llm4code/PengCHYR25@11CWEval ; fu2024constraineddecodingsecurecode . For instance, BaxBench DBLP:journals/corr/abs-2502-11844@12BaxBench designs correctness and security tests that are agnostic to frameworks and programming languages, but this generality makes it difficult to fully validate generated code, leading to under-reporting of vulnerabilities. To overcome these challenges, A.S.E emphasizes reproducibility and precision through customized vulnerability detection. For each data instance—corresponding to a specific project and CVE—A.S.E designs tailored static analysis rules that explicitly model the relevant sources, sinks, and taint propagation paths. These rules are calibrated to accurately detect the original vulnerability within the unmodified repository, ensuring that the evaluation framework can reliably assess whether regenerated code reintroduces or mitigates the same flaw. By automating this process at the project level, A.S.E achieves more reliable, scalable, and reproducible security assessments compared to prior approaches.
@@ 3.1 Design Philosophy
We aim to create a repository-level evaluation benchmark that mirrors real-world AI programming scenarios, offering a reliable framework for assessing the security of AI-generated code. To ensure accurate results, A.S.E. focuses on realistic data sources, task settings, and code assessment methods. Specifically, the core features of the A.S.E benchmark include the following three components:
@@
(ii) Task Settings: Real-world AI Programming Task Settings. To simulate realistic usage, A.S.E replicates AI programming assistants like Cursor by extracting code contexts—both intra-file and cross-file—directly from the repository. These contexts are then provided to LLMs for code generation, closely mimicking real-world AI programming scenarios. Moreover, the generated code is output in the form of diff files, allowing patches to be applied directly to the repository, further reflecting real software development practices.
@@
Following these guidelines, we introduce the A.S.E benchmark to evaluate in repository-level code generation regarding security. Next, we introduce the three key steps: the construction of the benchmark, the design of the code generation task, and the evaluation of the results.
@@
Guided by our design philosophy, we construct the A.S.E benchmark as shown in Figure 1 (a) and 2. To ensure realistic scenarios and relevant security expertise, we organize a dedicated team of ten contributors with strong backgrounds in cybersecurity and web development from top-tier universities, including five Ph.D. candidates and five master-level students. All contributors have extensive hands-on experience in vulnerability discovery, analysis, and remediation, with emphasis on common web vulnerabilities such as XSS, SQL injection, and path traversal. Each contributor adheres to strict secure coding standards and is familiar with static code analysis techniques. The construction of the benchmark proceeds in four stages: determining data sources, filtering candidate repository, expert-guided refinement and quality filtering, and dataset expansion. The following subsections will provide a detailed explanation of each step.
Step 1: Determining Data Sources. We collect real-world software projects from both enterprise-internal vulnerability repositories and public platforms like GitHub. This stage ensures representativeness and diversity, so that the dataset covers a broad range of software types and vulnerability scenarios.
Step 2: Filtering candidate repository. Then we apply strict criteria to filter candidate repositories, ensuring both project quality and vulnerability completeness. Only repositories with active maintenance, sufficient test coverage, and consistent development practices are retained. In addition, each repository must contain at least one publicly disclosed CVE accompanied by a complete patch diff at the commit level. For every vulnerability case, detailed metadata—such as a textual description, CVE identifier, severity rating, and reproducible steps—must be available. These conditions guarantee that the selected projects provide realistic software contexts and verifiable vulnerability information, thereby ensuring high data quality and reproducibility for subsequent stages.
@@
Step 3: Expert-Guided Refinement and Quality Control. To ensure that the collected data is security-relevant and represents real-world vulnerabilities that are both detectable and reproducible, we refine the dataset further through expert-guided analysis and systematic filtering. After the initial screening, security experts examine each candidate repository to precisely annotate vulnerable code regions. The system then extracts surrounding context—including function signatures, API definitions, and call chains—to reconstruct the code environment. For every vulnerability, we design CodeQL or Joern queries and verify the corresponding taint propagation paths. Projects with ambiguous or inconsistent results, or with code that current tools cannot analyze with sufficient precision, are discarded. For the retained cases, detection rules are customized and refined to improve accuracy and ensure alignment with confirmed vulnerabilities. Once validated, the labeled vulnerable code is removed to create a fill-in-the-code setting. By combining the functional description of the vulnerability with its extracted context, we generate structured prompts that allow models to reason over repository-level structures and logic rather than isolated snippets. After expert review, 40 repositories with verified CVE records are selected as the seed dataset, each with a baseline commit serving as a stable starting point for task construction and evaluation. This process guarantees authenticity, reliability, and reproducibility across all benchmark tasks.
@@
Step 4: Dataset Expansion. After constructing base tasks from real-world vulnerability repositories, we expand the dataset to increase both volume and coverage under a semantics-preserving constraint. Two categories of transformations are applied. The first is semantic transformations, which diversify surface expressions by systematically renaming variables and functions or substituting equivalent APIs. The second is structural transformations, which alter control flow, refactor call graphs, or reorganize file layouts to introduce structural variation. These operations preserve functional behavior and code semantics while modifying implementation details, thereby reducing overlap with public repository code that may appear in model training corpora and mitigating potential data contamination. The expanded variants enable a more comprehensive assessment of model robustness and generalization.
Figure 3: Statistics of A.S.E benchmark.
3.2.1 General Statistics
After the construction, we obtain the A.S.E. benchmark, which comprises $120$ repository-level vulnerability instances, including $40$ seed repositories and $80$ mutated variants.
Specifically, the benchmark targets four vulnerability categories that are common in real web projects. Each category maps one-to-one to a corresponding CWE entry: SQL Injection ($29.2\%$, CWE-89), Path Traversal ($26.7\%$, CWE-22), Cross-Site Scripting ($25.0\%$, CWE-79), and Command Injection ($19.2\%$, CWE-78). This mapping defines the dataset at the CWE level and limits it to these four categories.
@@ 3.2.1 General Statistics
From the perspective of programming languages, A.S.E encompasses five mainstream environments that reflect realistic multi-language software development. The distribution is concentrated in PHP ($50.0\%$), followed by Python ($19.2\%$), Go ($14.2\%$), JavaScript ($14.2\%$), and Java ($2.5\%$). This distribution highlights the prevalence of vulnerable PHP code in real-world applications while also offering opportunities to explore cross-language generalization in evaluation.
3.3 Model Code Generation

### v3

@@ Abstract
The increasing adoption of large language models (LLMs) in software engineering necessitates rigorous security evaluation of their generated code. However, existing benchmarks often lack relevance to real-world AI-assisted programming scenarios, making them inadequate for assessing the practical security risks associated with AI-generated code in production environments. To address this gap, we introduce A.S.E (AI Code Generation Security Evaluation), a repository-level evaluation benchmark designed to closely mirror real-world AI programming tasks, offering a comprehensive and reliable framework for assessing the security of AI-generated code. Our evaluation of leading LLMs on A.S.E reveals several key findings. In particular, current LLMs still struggle with secure coding. The complexity in repository-level scenarios presents challenges for LLMs that typically perform well on snippet-level tasks. Moreover, a larger reasoning budget does not necessarily lead to better code generation. These observations offer valuable insights into the current state of AI code generation and help developers identify the most suitable models for practical tasks. They also lay the groundwork for refining LLMs to generate secure and efficient code in real-world applications.
@@ 1 Introduction
The rapid advancement of large language models (LLMs) has greatly enhanced the AI programming ecosystem, with tools like Cursor @Cursor and Claude Code @claude-code enabling developers to choose models that best fit their tasks. These AI assistants significantly improve programming efficiency, leading to a surge of AI-generated code in production environments. However, research @8SecurityEval ; DBLP:journals/corr/abs-2502-11844@12BaxBench ; DBLP:conf/llm4code/PengCHYR25@11CWEval ; DBLP:conf/satml/HajipourHHSF24@9CodeLMSec ; DBLP:journals/corr/abs-2506-05692@13SafeGenBench ; fu2024constraineddecodingsecurecode ; he2023large ; pearce2025asleep ; wang2024your has shown that such code can harbor security vulnerabilities, posing serious risks such as data breaches or system failures pearce2025asleep ; @8SecurityEval ; fu2023security ; li2017large . Relying on developers to ensure the security of AI-generated code can be highly challenging given the complexity of modern software systems. Therefore, there is a pressing need to identify and utilize AI models that are capable of generating secure code.
@@
Despite the numerous benchmarks DBLP:conf/nips/HendrycksBKMAGB21@21APPS ; dou2024stepcoder ; DBLP:journals/corr/abs-2107-03374@17HumanEval ; DBLP:journals/corr/abs-2108-07732@19MBPP ; DBLP:journals/corr/abs-2506-05692@13SafeGenBench ; DBLP:journals/corr/abs-2502-11844@12BaxBench ; DBLP:conf/llm4code/PengCHYR25@11CWEval ; DBLP:journals/corr/abs-2405-00218@10constrained ; DBLP:conf/satml/HajipourHHSF24@9CodeLMSec ; @8SecurityEval ; wang2024your developed by both academia and industry to evaluate AI-generated code, most of them  DBLP:conf/nips/HendrycksBKMAGB21@21APPS ; dou2024stepcoder ; DBLP:journals/corr/abs-2107-03374@17HumanEval ; DBLP:journals/corr/abs-2108-07732@19MBPP primarily focus on code quality, such as syntax correctness and functional accuracy, while overlooking critical security considerations. Although some benchmarks DBLP:journals/corr/abs-2506-05692@13SafeGenBench ; DBLP:journals/corr/abs-2502-11844@12BaxBench ; DBLP:conf/llm4code/PengCHYR25@11CWEval ; DBLP:journals/corr/abs-2405-00218@10constrained ; DBLP:conf/satml/HajipourHHSF24@9CodeLMSec ; @8SecurityEval ; wang2024your attempt to address code security (as shown in Table 1), they are often inadequate to assess the actual security risks of AI-generated code in real-world production scenarios due to several key reasons: (a) Limited relevance to real-world data. These datasets are typically sourced from human-curated synthetic code snippets, which have limited relevance to the functional and security scenarios of real-world projects. (b) Code generation tasks detached from real-world AI programming. Their code generation tasks are typically limited to isolated code snippets, focusing solely on functional descriptions without considering the context within files or projects, which does not align with mainstream AI programming paradigm. (c) Unreliable code evaluation methods. Security assessments of generated code often rely on manual or LLM-based judgment, which are unreliable and difficult to automate or reproduce consistently. This gap poses a significant challenge for both developers and organizations seeking to integrate AI-generated code securely into their systems.
@@
Building on these design principles, the A.S.E benchmark includes 120 repository-level instances, consisting of 40 seed dataset collected from GitHub and 80 mutated variants generated through semantic and structural mutation techniques, such as identifier renaming and control-flow reshaping. These variants are introduced to mitigate data leakage risks, ensuring that the evaluation reflects the LLM’s capabilities rather than its memorization. To closely simulate real-world code generation processes, A.S.E uses the BM25 algorithm DBLP:journals/ftir/RobertsonZ09@bm25 to automatically extract relevant code context from the repository when designing code generation tasks. These code generation tasks cover four common vulnerability types in real-world web projects. We focus on the web domain due to the primary application of AI programming in web development in production environments. These vulnerability-related codes to be generated are closely tied to the projects’ business logic, requiring models to understand both project code and security considerations. A.S.E also incorporates five commonly used web development languages to assess how programming languages influence model performance. Furthermore, A.S.E integrates customized static application security testing (SAST) rules for security evaluation on each data instance in the dataset. While A.S.E aims to evaluate the security of AI-generated code, it also considers code quality and stability. Given that security is meaningful only when code correctness is ensured, and in light of potential hallucinations in large model outputs, A.S.E uses multi-dimensional metrics that assess code security, quality, and generation stability for a comprehensive evaluation of the model’s capabilities.
@@
Based on A.S.E, we evaluated 26 mainstream commercial or open-source models under the same experimental setup, leading to several key findings. First, existing LLMs still face significant challenges in secure coding. All models fall short in terms of security performance compared to their code quality performance (such as syntax correctness). Even the best-performing model, Claude-3.7-Sonnet, achieved only a total score of 52.79 in our evaluation. Second, A.S.E introduces significant complexity in repository-level scenarios, which presents a challenge for LLMs that typically perform well on snippet-level tasks. For example, although GPT-3 excels on SafeGenBench DBLP:journals/corr/abs-2506-05692@13SafeGenBench , its performance on the A.S.E. benchmark drops, falling behind many other models. Third, slow-thinking configurations, which allocate more deliberate computation or multi-step reflection, tend to underperform compared to fast-thinking configurations that rely on concise, direct decoding. This suggests that a larger reasoning budget does not necessarily lead to better code generation. These observations offer valuable insights into the current state of AI code generation, helping developers select the most appropriate models for their specific tasks. Furthermore, they provide a foundation for refining LLMs, enhancing their ability to generate secure and efficient code in real-world applications.
@@ 2 Related Work
Table 1: Comparison of Security-Oriented Code Generation Evaluation Datasets.
@@
SafeGenBench Snippet Human-Curated Synthetic Simplified Programming Tasks SAST + LLM
@@
CWEval Snippet Human-Curated Synthetic Simplified Programming Tasks Test Cases
CODEGUARD+ Snippet Human-Curated Synthetic Simplified Programming Tasks Test Cases
CodeLMSec Snippet Human-Curated Synthetic Simplified Programming Tasks SAST
SecurityEval Snippet Human-Curated Synthetic Simplified Programming Tasks SAST + Manual
@@ 2.1 Relevance to Real-world Scenarios
2.2 Code Assessment Methods
Prior security-oriented benchmarks have adopted diverse methods for assessing the security of AI-generated code, yet each comes with notable limitations. (a) Manual evaluation offers flexibility but is labor-intensive and difficult to scale. For example, SecurityEval @8SecurityEval relies on expert validation, which makes it difficult to automate or scale for evaluating new models. (b) LLM-as-judge approaches scale more easily and capture semantic nuances, but their judgments are highly sensitive to prompt design, model version, and decoding randomness, making results unreliable and hard to reproduce (DBLP:journals/corr/abs-2312-04724@25CyberSecEval, ). (c) Generic static analysis (SAST) tools provide deterministic rule-based detection, yet often suffer from false positives and false negatives when applied across different languages or without contextual calibration DBLP:conf/satml/HajipourHHSF24@9CodeLMSec ; DBLP:journals/corr/abs-2506-05692@13SafeGenBench . (d) Test-case-based assessments are straightforward but inherently limited DBLP:journals/corr/abs-2502-11844@12BaxBench ; DBLP:conf/llm4code/PengCHYR25@11CWEval ; fu2024constraineddecodingsecurecode . For instance, BaxBench DBLP:journals/corr/abs-2502-11844@12BaxBench designs correctness and security tests that are agnostic to frameworks and programming languages, but this generality makes it difficult to fully validate generated code, leading to under-reporting of vulnerabilities. To overcome these challenges, A.S.E emphasizes reproducibility and precision through customized vulnerability detection. For each data instance corresponding to a specific project and CVE, A.S.E designs tailored static analysis rules that explicitly model the relevant sources, sinks, and taint propagation paths. These rules are calibrated to accurately detect the original vulnerability within the unmodified repository, ensuring that the evaluation framework can reliably assess whether regenerated code reintroduces or mitigates the same flaw. By automating this process at the project level, A.S.E achieves more reliable, scalable, and reproducible security assessments compared to prior approaches.
@@ 3.1 Design Philosophy
We aim to create a repository-level evaluation benchmark that mirrors real-world AI programming scenarios, offering a reliable framework for assessing the security of AI-generated code. To ensure accurate results, A.S.E. focuses on realistic data sources, task settings, and code assessment methods. Specifically, the core features of the A.S.E benchmark are designed around the following principles:
@@
(ii) Task Settings: Practical Simulations of AI Programming Workflows. To simulate realistic usage, A.S.E replicates AI programming assistants like Cursor by extracting code contexts—both intra-file and cross-file—directly from the repository. These contexts are then provided to LLMs for code generation, closely mimicking real-world AI programming scenarios. Moreover, the generated code is output in the form of diff files, allowing patches to be applied directly to the repository, further reflecting real software development practices.
@@
Following these guidelines, we introduce the A.S.E benchmark to evaluate in repository-level code generation regarding security. After that, we detail the three key steps: benchmark construction, task design, and result evaluation.
@@
Guided by our design philosophy, we construct the A.S.E benchmark as shown in Figure 1 (a) and Figure 2. To ensure realistic scenarios and relevant security expertise, we organize a dedicated team of ten contributors with strong backgrounds in cybersecurity and web development from top-tier universities, including five Ph.D. candidates and five master-level students. All contributors have extensive hands-on experience in vulnerability discovery, analysis, and remediation, with emphasis on common web vulnerabilities such as XSS, SQL injection, and path traversal. Each contributor adheres to strict secure coding standards and is familiar with static code analysis techniques. The construction of the benchmark proceeds in four stages: determining data sources, filtering candidate repository, expert-guided refinement and quality filtering, and dataset expansion. The following subsections will provide a detailed explanation of each step.
Step 1: Determining Data Sources. To ensure that our code generation tasks are both security-sensitive and representative of real-world scenarios, we begin by collecting source data directly from CVE vulnerabilities. Specifically, we gather CVE records and their corresponding repository information from both public vulnerability databases and enterprise-internal repositories. For each repository, we additionally require accessible commit history, which allows us to accurately locate the vulnerable code and later design code generation tasks. As a result of this initial step, we obtain more than $100,000$ raw CVE entries as the starting point for benchmark construction.
Step 2: Filtering Candidate Repositories. After collecting raw CVE entries, we perform a strict filtering process to ensure both project quality and vulnerability completeness. Specifically, we first narrow the dataset by retaining only entries that (i) map to web-relevant categories in the 2024 Top CWE list CWE_Top25_2024 and (ii) provide traceable vulnerability fix commit contexts. This restriction reflects our focus on web applications, which constitute one of the most common real-world domains for AI programming, and ensures that every vulnerability can be grounded in a concrete, analyzable code change rather than an abstract description. Then we further filter the repositories by requiring either active monthly maintenance or a popularity threshold of more than $1,000$ GitHub stars. This step eliminates abandoned or low-quality projects and ensures that the retained repositories exhibit sufficient code complexity and practical relevance, thereby making the benchmark more representative of real-world software engineering environments. After applying these criteria, the dataset is reduced to approximately $50,000$ candidate repositories.
Furthermore, we apply SAST tools (e.g., CodeQL CodeQL and Joern Yamaguchi2014ModelingAD ) to these repositories and intersect reported findings with the lines modified in the corresponding commits. By retaining only cases where SAST detections coincide with the lines modified in the corresponding commits, we simultaneously reduce false positives and enforce a verifiable vulnerability–fix causal chain. This design ensures that each CVE instance is (i) detectable by automated static analysis, confirming its practical observability, and (ii) accurately linked to the correct fixing commit, guaranteeing that the patched code genuinely corresponds to the target vulnerability. These conditions provide reliable vulnerability evidence that can support subsequent expert-guided analysis. As a result, this stage yields 199 high-confidence candidates.
@@
Step 3: Expert-Guided Refinement and Quality Control. To ensure the final benchmark is genuinely security-relevant and reflects real-world vulnerabilities that are both detectable and reproducible, we refine the dataset through expert annotation and validation. Specifically, we first conduct an initial manual review to further control data quality. At this stage, security experts remove obvious false positives introduced by static analysis tools and discard commits that modify an excessive number of files (e.g., more than 10), since large-scale changes obscure vulnerability localization and hinder precise labeling. Building on the cleaned candidate set, we then proceed with a fine-grained expert analysis that focuses on the vulnerabilities themselves. Security experts precisely annotate the vulnerable code regions, reconstruct the relevant execution context (e.g., source/sink signatures, API definitions, call chains), and design targeted CodeQL/Joern queries to validate taint propagation paths. Once validated, the labeled vulnerable code is removed to create a fill-in-the-code setting. By combining the functional description of the vulnerability with its extracted context, we generate structured prompts that require models to reason over repository-level structures and logic rather than isolated snippets. After final expert review, 40 repositories with verified CVE records are retained as the seed dataset, each anchored at a baseline commit that provides a stable starting point for task construction and evaluation. This expert-driven process guarantees authenticity, reliability, and reproducibility across all benchmark tasks.
@@
Step 4: Dataset Expansion. After constructing base tasks from real-world CVE vulnerabilities and corresponding repositories, we expand the dataset to increase both volume and coverage under a semantics-preserving constraint. Two categories of transformations are applied. The first is semantic transformations, which diversify surface expressions by systematically renaming variables and functions or substituting equivalent APIs. The second is structural transformations, which alter control flow, refactor call graphs, or reorganize file layouts to introduce structural variation. These operations preserve functional behavior and code semantics while modifying implementation details, thereby reducing overlap with public repository code that may appear in model training corpora and mitigating potential data contamination. The expanded variants enable a more comprehensive assessment of model robustness and generalization. Through these transformations, we generate 80 additional variants from the 40 seed repositories, yielding a total of 120 benchmark instances.
Figure 3: Benchmark Construction Funnel.
General Statistics. Figure 3 illustrates the data reduction pipeline from the initial collection of raw CVE entries to the final benchmark. The funnel chart presents the number of instances retained after each stage of filtering and refinement, showing how the dataset was progressively narrowed from a large pool of raw vulnerabilities to a carefully curated benchmark. The resulting A.S.E benchmark comprises $120$ repository-level vulnerability instances and the overall composition is illustrated in Figure 4.
Figure 4: Statistics of A.S.E benchmark, including the distribution of top CWE categories, programming languages, and repositories.
Specifically, the dataset targets four categories of vulnerabilities that are widely prevalent in real-world web projects, each aligned with a CWE entry: SQL Injection ($29.2\%$, CWE-89), Path Traversal ($26.7\%$, CWE-22), Cross-Site Scripting ($25.0\%$, CWE-79), and Command Injection ($19.2\%$, CWE-78). This mapping defines the dataset at the CWE level, ensuring that evaluation tasks align with security-critical issues that LLMs must account for when generating code in real-world development. Each category captures a distinct challenge where secure functionality requires the model not only to implement business logic correctly but also to avoid unsafe coding patterns:
  • Cross-Site Scripting (XSS): evaluates whether the model can generate web logic (e.g., input/output rendering) while preventing injection of malicious scripts into trusted contexts.
  • SQL Injection (SQLI): tests whether the model, when generating database operation logic, properly handles user input and avoids unsafe SQL statement construction.
  • Path Traversal: examines if the model can implement file access functionality without exposing sensitive paths outside the intended directory scope.
  • Command Injection: assesses whether the model can generate code involving system interactions while preventing the execution of unauthorized operating system commands.
@@
From a language perspective, A.S.E spans five mainstream programming environments to reflect realistic multi-language software development. The distribution is concentrated in PHP ($50.0\%$), followed by Python ($19.2\%$), Go ($14.2\%$), JavaScript ($14.2\%$), and Java ($2.5\%$). This distribution highlights the dominance of PHP in vulnerability-prone web applications while also enabling evaluation of model generalization across diverse programming languages.
We also analyze the size of the vulnerable code that define each code generation task. Specifically, the number of vulnerable lines of code (LOC) per task varies substantially, with an average of 35.77, a median of 18, and a range of [2–415]. These statistics characterize the functional code fragments that models are required to regenerate, highlighting the variation in task complexity—from small, localized code edits spanning only a few lines to larger segments involving multiple statements or function bodies. This diversity ensures that the benchmark captures both simple and complex generation scenarios under realistic repository-level settings.
For tooling integration, we incorporate two state-of-the-art static analysis frameworks—CodeQL and Joern—which are packaged into containerized environments to ensure reproducibility and ease of deployment. Each tool is applied to 50% of the benchmark instances, providing complementary static analysis capabilities for security evaluation. The containerization not only standardizes the execution environment across different platforms but also guarantees that results are consistent and reproducible.
3.3 Code Generation